how many moles of PCl5 can be produced from 58.0 g of Cl2 (excess P4)


1
Expert's answer
2022-04-08T03:51:03-0400

m(Cl2) = 58.0 g;

M(Cl2) = 71 g/mol;

n(Cl2) = m(Cl2)/M(Cl2) = 58/71 = 0.817 mol;

P4 + 10Cl2 = 4PCl5;

Since P4 is the excess reagent, then by the chemical reaction:

n(PCl5) = 4/10 * n(Cl2) = 4/10 * 0.817 = 0.327 mol.

Answer: 0.327 mol.